# A splitting, discontinuous Galerkin solver for the cell-by-cell electroneutral Nernst-Planck framework

### v1

@@ 2.1 Governing equations
$\frac{\partial c_{k}}{\partial t}+\operatorname{\nabla\cdot}\mathbf{J}_{k}$ $=0$ in $\qquad\text{in }\Omega,$ (1)
$F\sum_{k\in K}z_{k}\operatorname{\nabla\cdot}\mathbf{J}_{k}$ $=0$ in $\qquad\text{in }\Omega,$ (2)
@@
$\mathbf{J}_{k}=-D_{k}\nabla c_{k}-c_{k}z_{k}D_{k}\psi\nabla\phi,\qquad\text{in}\Omega$ (8)
@@
$\alpha_{k}=\frac{D_{k}z_{k}^{2}c_{k}}{\sum_{l\in K}D_{l}z_{l}^{2}c_{l}}\qquad\text{ on }\Gamma.$ (9)
@@
$I_{\text{ch},k}=g_{k}(\phi_{M}-E_{k}),\qquad E_{k}=\frac{RT}{z_{k}F}\ln{\frac{c_{k,e}}{c_{k,i}}}\qquad\text{ on }\Gamma,$ (10)
@@
$\frac{\partial\phi_{M}}{\partial t}=\frac{1}{C_{M}}(I_{M}-I_{\text{ch}})\qquad\text{ on }\Gamma,$ (11)
$\frac{\partial\mathbf{s}}{\partial t}=F(\phi_{M},\mathbf{s})\qquad\text{ on }\Gamma.$ (12)
@@ 2.2 Formulation with fewer concentrations
$0=\sum_{k=1}^{K}z_{k}c_{k}\qquad\text{in }\Omega.$ (13)
@@
$c_{m}=-\frac{1}{z_{m}}\sum_{k=1}^{m-1}z_{k}c_{k}\qquad\text{in }\Omega.$ (14)
@@ 2.3 Boundary and initial conditions
$\begin{split}c_{k}(\mathbf{x},0)&=c_{k}^{0}=\begin{cases}c_{k,i}^{0}&\quad\text{if }\mathbf{x}\in\Omega_{i}\\ c_{k,e}^{0}&\quad\text{if }\mathbf{x}\in\Omega_{e},\end{cases}\end{split}$ (15)
$\phi_{M}(\mathbf{x},0)$ $=\phi_{M}^{0}\qquad\qquad\quad\quad\mathbf{x}\in\Gamma,$ (16)
@@
$0=\sum_{k=1}^{K}z_{k}c_{k}^{0}\qquad\text{in }\Omega.$ (17)
@@
$\mathbf{J}_{k}(\mathbf{x},t)\cdot\mathbf{n}=0\qquad\mathbf{x}\in\partial\Omega.$ (18)
@@ 3.1 Temporal discretization and splitting scheme
$\frac{\partial c_{k}}{\partial t}\approx\frac{c_{k}^{n}-c_{k}^{n-1}}{\Delta t}\quad\text{and}\quad\frac{\partial\phi_{M}}{\partial t}\approx\frac{\phi_{M}^{n}-\phi_{M}^{n-1}}{\Delta t}$ (19)
@@
    $F\sum_{k\in K}z_{k}\operatorname{\nabla\cdot}{\mathbf{J}_{k}^{n}}$ $=0,$ (20)
    ${\mathbf{J}_{k}^{n}}$ $=-D_{k}\nabla c_{k}^{n-1}-z_{k}D_{k}\psi c_{k}^{n-1}\nabla\phi^{n}.$ (21)
@@
    ${\Delta t}^{-1}c_{k}^{n}+\operatorname{\nabla\cdot}\mathbf{J}_{k}^{n}$ $={\Delta t}^{-1}c_{k}^{n-1},$ (22)
    $\mathbf{J}_{k}^{n}$ $=-D_{k}\nabla c_{k}^{n}-z_{k}D_{k}\psi c_{k}^{n}\nabla\phi^{n}$ (23)
@@
$I_{M}=\frac{C_{M}}{\Delta t}(\phi_{i}^{n}-\phi_{e}^{n}-f)\qquad\text{ on }\Gamma,$ (24)
@@
$f=f^{n-1}=\phi_{M}^{n-1}-\frac{\Delta t}{C_{M}}I_{\text{ch}}.$ (25)
@@
$\mathbf{J}_{k,i}\cdot\mathbf{n}_{i}=\frac{\alpha_{k,i}C_{M}}{Fz_{k}\Delta t}(\phi_{i}^{n}-\phi_{e}^{n}-g_{k,i})\qquad\text{ on }\Gamma,$ (26)
$-\mathbf{J}_{k,e}\cdot\mathbf{n}_{e}=\frac{\alpha_{k,e}C_{M}}{Fz_{k}\Delta t}(\phi_{i}^{n}-\phi_{e}^{n}-g_{k,e})\qquad\text{ on }\Gamma,$ (27)
@@
$g_{k}=g_{k}^{n-1}=\phi_{M}^{n-1}-\frac{\Delta t}{C_{M}\alpha_{k}}.$ (28)
@@ 3.2.1 Continuous problems
$\frac{1}{\Delta t}\int_{\Omega}c_{k}^{n}v_{k}+a_{c_{k}}(c_{k}^{n},v_{k})=\frac{1}{\Delta t}\int_{\Omega}c_{k}^{n-1}v_{k}+l_{c_{k}}(v_{k})\quad\forall v_{k}\in V,$ (31)
@@ 3.2.2 DG discretization
$[u]=u_{1}-u_{2},\quad\{u\}=\frac{u_{1}+u_{2}}{2}\quad\text{on }e\in\mathcal{F}_{i}^{h}\cup\mathcal{F}^{h}_{m},$ (32)
@@
$\sum_{E\in\mathcal{T}_{h}}\frac{1}{\Delta t}\int_{E}c_{k}^{n}v_{k}+a_{c_{k}}^{h}(c_{k}^{n},v_{k})=\sum_{E\in\mathcal{T}_{h}}\frac{1}{\Delta t}\int_{E}c_{k}^{n-1}v_{k}+l_{c_{k}}^{h}(v_{k})\quad\forall v_{k}\in V^{h}_{p},$ (34)
@@ 3.3 Extension of scheme for active (ODE governed) membrane models
$g_{k}=g_{k}^{n-1}=\phi_{M}^{n-1}-\frac{\Delta t}{C_{M}\alpha_{k}}I_{\text{ch},k}+\frac{\Delta t}{C_{M}}I_{\text{ch}}$ (36)
@@ 5.1 Model A: Smooth manufactured solutions
$c_{\textrm{Na},i}$ $=0.7+0.3\sin(2\pi x)\sin(2\pi y)$ in $\qquad\text{in }\Omega_{i},$ (37)
@@
$c_{\textrm{Na},i}$ $=1+x+y+0.3\cos(2\pi t)$ in $\qquad\text{in }\Omega_{i},$ (38)
@@ 5.2 Model B: Idealized 2D axon with active membrane model
$I_{\rm syn}(\mathbf{x},t)=g_{\rm syn}f(\mathbf{x})e^{-\frac{t}{\tau}}(\phi_{M}-E_{\rm Na}),$ (39)
@@
$f(\mathbf{x})=\begin{cases}1&\text{if }x\leq 1\mu\rm m\\ 0&\text{else}\end{cases}\qquad\mathbf{x}=(x,y)\in\Gamma.$ (40)
@@ 5.3 Model C: Idealized 3D axons with active membrane model
$f(\mathbf{x})=\begin{cases}1&\text{if }x\leq 20\mu\rm m\\ 0&\text{else}\end{cases}\qquad\mathbf{x}=(x,y,z)\in\Gamma.$ (41)
@@ 5.4 Model D: Morphologically realistic neuron with spatially varying membrane model
$f(\mathbf{x})=\begin{cases}1&\text{if }x\leq-125\mu{\rm m},y\geq 140\mu{\rm m},z\leq-80\mu{\rm m}\\ 0&\text{else}\end{cases}\quad\mathbf{x}=(x,y,z)\in\Gamma.$ (42)
@@ 5.6 Peclet number
$\textrm{Pe}=\frac{Lz_{k}\frac{FD}{RT}\nabla\phi}{D}=\frac{Lz_{k}F\nabla\phi}{RT},$ (44)
@@
$\frac{|D\nabla c_{k}|-|z_{k}c\frac{FD}{RT}\nabla\phi|}{\rm{max}(|D\nabla c_{k}|,\epsilon)}.$ (45)
@@ 6.7 Properties of the DG scheme in realistic brain tissue geometries
[…]
@@
[…]
@@ Acknowledgments
Ada J. Ellingsrud acknowledges support from the Research Council of Norway via FRIPRO grant #324239 (EMIx) and from the national infrastructure for computational science in Norway, Sigma2, via grant #NN8049K. Miroslav Kuchta acknowledges support from the Research Council of Norway via DataSim #303362. We thank Rami Masri for valuable discussions and input on discontinuous Galerkin methods, Pietro Benedusi for running the parallel scaling experiments and Jørgen Dokken for his contributions in setting up the associated software repository.

### v2

@@ 2.1 Governing equations
(1) $\frac{\partial c_{k}}{\partial t}+\operatorname{\nabla\cdot}\mathbf{J}_{k}$ $=0$ in $\qquad\text{in }\Omega,$
(2) $F\sum_{k\in K}z_{k}\operatorname{\nabla\cdot}\mathbf{J}_{k}$ $=0$ in $\qquad\text{in }\Omega,$
@@
(8) $\mathbf{J}_{k}=-D_{k}\nabla c_{k}-c_{k}z_{k}D_{k}\psi\nabla\phi,\qquad\text{in}\Omega$
@@
(9) $\alpha_{k}=\frac{D_{k}z_{k}^{2}c_{k}}{\sum_{l\in K}D_{l}z_{l}^{2}c_{l}}\qquad\text{ on }\Gamma.$
@@
(10) $I_{\text{ch},k}=g_{k}(\phi_{M}-E_{k}),\qquad E_{k}=\frac{RT}{z_{k}F}\ln{\frac{c_{k,e}}{c_{k,i}}}\qquad\text{ on }\Gamma,$
@@
(11) $\frac{\partial\phi_{M}}{\partial t}=\frac{1}{C_{M}}(I_{M}-I_{\text{ch}})\qquad\text{ on }\Gamma,$
(12) $\frac{\partial\mathbf{s}}{\partial t}=F(\phi_{M},\mathbf{s})\qquad\text{ on }\Gamma.$
@@ 2.2 Formulation with fewer concentrations
(13) $0=\sum_{k=1}^{K}z_{k}c_{k}\qquad\text{in }\Omega.$
@@
(14) $c_{m}=-\frac{1}{z_{m}}\sum_{k=1}^{m-1}z_{k}c_{k}\qquad\text{in }\Omega.$
@@ 2.3 Boundary and initial conditions
(15) $\begin{split}c_{k}(\mathbf{x},0)&=c_{k}^{0}=\begin{cases}c_{k,i}^{0}&\quad\text{if }\mathbf{x}\in\Omega_{i}\\ c_{k,e}^{0}&\quad\text{if }\mathbf{x}\in\Omega_{e},\end{cases}\end{split}$
(16) $\phi_{M}(\mathbf{x},0)$ $=\phi_{M}^{0}\qquad\qquad\quad\quad\mathbf{x}\in\Gamma,$
@@
(17) $0=\sum_{k=1}^{K}z_{k}c_{k}^{0}\qquad\text{in }\Omega.$
@@
(18) $\mathbf{J}_{k}(\mathbf{x},t)\cdot\mathbf{n}=0\qquad\mathbf{x}\in\partial\Omega.$
@@ 3.1 Temporal discretization and splitting scheme
(19) $\frac{\partial c_{k}}{\partial t}\approx\frac{c_{k}^{n}-c_{k}^{n-1}}{\Delta t}\quad\text{and}\quad\frac{\partial\phi_{M}}{\partial t}\approx\frac{\phi_{M}^{n}-\phi_{M}^{n-1}}{\Delta t}$
@@
    (20) $F\sum_{k\in K}z_{k}\operatorname{\nabla\cdot}{\mathbf{J}_{k}^{n}}$ $=0,$
    (21) ${\mathbf{J}_{k}^{n}}$ $=-D_{k}\nabla c_{k}^{n-1}-z_{k}D_{k}\psi c_{k}^{n-1}\nabla\phi^{n}.$
@@
    (22) ${\Delta t}^{-1}c_{k}^{n}+\operatorname{\nabla\cdot}\mathbf{J}_{k}^{n}$ $={\Delta t}^{-1}c_{k}^{n-1},$
    (23) $\mathbf{J}_{k}^{n}$ $=-D_{k}\nabla c_{k}^{n}-z_{k}D_{k}\psi c_{k}^{n}\nabla\phi^{n}$
@@
(24) $I_{M}=\frac{C_{M}}{\Delta t}(\phi_{i}^{n}-\phi_{e}^{n}-f)\qquad\text{ on }\Gamma,$
@@
(25) $f=f^{n-1}=\phi_{M}^{n-1}-\frac{\Delta t}{C_{M}}I_{\text{ch}}.$
@@
(26) $\mathbf{J}_{k,i}\cdot\mathbf{n}_{i}=\frac{\alpha_{k,i}C_{M}}{Fz_{k}\Delta t}(\phi_{i}^{n}-\phi_{e}^{n}-g_{k,i})\qquad\text{ on }\Gamma,$
(27) $-\mathbf{J}_{k,e}\cdot\mathbf{n}_{e}=\frac{\alpha_{k,e}C_{M}}{Fz_{k}\Delta t}(\phi_{i}^{n}-\phi_{e}^{n}-g_{k,e})\qquad\text{ on }\Gamma,$
@@
(28) $g_{k}=g_{k}^{n-1}=\phi_{M}^{n-1}-\frac{\Delta t}{C_{M}\alpha_{k}}.$
@@ 3.2.1 Continuous problems
(31) $\frac{1}{\Delta t}\int_{\Omega}c_{k}^{n}v_{k}+a_{c_{k}}(c_{k}^{n},v_{k})=\frac{1}{\Delta t}\int_{\Omega}c_{k}^{n-1}v_{k}+l_{c_{k}}(v_{k})\quad\forall v_{k}\in V,$
@@ 3.2.2 DG discretization
(32) $[u]=u_{1}-u_{2},\quad\{u\}=\frac{u_{1}+u_{2}}{2}\quad\text{on }e\in\mathcal{F}_{i}^{h}\cup\mathcal{F}^{h}_{m},$
@@
(34) $\sum_{E\in\mathcal{T}_{h}}\frac{1}{\Delta t}\int_{E}c_{k}^{n}v_{k}+a_{c_{k}}^{h}(c_{k}^{n},v_{k})=\sum_{E\in\mathcal{T}_{h}}\frac{1}{\Delta t}\int_{E}c_{k}^{n-1}v_{k}+l_{c_{k}}^{h}(v_{k})\quad\forall v_{k}\in V^{h}_{p},$
@@ 3.3 Extension of scheme for active (ODE governed) membrane models
(36) $g_{k}=g_{k}^{n-1}=\phi_{M}^{n-1}-\frac{\Delta t}{C_{M}\alpha_{k}}I_{\text{ch},k}+\frac{\Delta t}{C_{M}}I_{\text{ch}}$
@@ 5.1 Model A: Smooth manufactured solutions
(37) $c_{\textrm{Na},i}$ $=0.7+0.3\sin(2\pi x)\sin(2\pi y)$ in $\qquad\text{in }\Omega_{i},$
@@
(38) $c_{\textrm{Na},i}$ $=1+x+y+0.3\cos(2\pi t)$ in $\qquad\text{in }\Omega_{i},$
@@ 5.2 Model B: Idealized 2D axon with active membrane model
(39) $I_{\rm syn}(\mathbf{x},t)=g_{\rm syn}f(\mathbf{x})e^{-\frac{t}{\tau}}(\phi_{M}-E_{\rm Na}),$
@@
(40) $f(\mathbf{x})=\begin{cases}1&\text{if }x\leq 1\mu\rm m\\ 0&\text{else}\end{cases}\qquad\mathbf{x}=(x,y)\in\Gamma.$
@@ 5.3 Model C: Idealized 3D axons with active membrane model
(41) $f(\mathbf{x})=\begin{cases}1&\text{if }x\leq 20\mu\rm m\\ 0&\text{else}\end{cases}\qquad\mathbf{x}=(x,y,z)\in\Gamma.$
@@ 5.4 Model D: Morphologically realistic neuron with spatially varying membrane model
(42) $f(\mathbf{x})=\begin{cases}1&\text{if }x\leq-125\mu{\rm m},y\geq 140\mu{\rm m},z\leq-80\mu{\rm m}\\ 0&\text{else}\end{cases}\quad\mathbf{x}=(x,y,z)\in\Gamma.$
@@ 5.6 Peclet number
(44) $\textrm{Pe}=\frac{Lz_{k}\frac{FD}{RT}\nabla\phi}{D}=\frac{Lz_{k}F\nabla\phi}{RT},$
@@
(45) $\frac{|D\nabla c_{k}|-|z_{k}c\frac{FD}{RT}\nabla\phi|}{\rm{max}(|D\nabla c_{k}|,\epsilon)}.$
@@ 6.7 Properties of the DG scheme in realistic brain tissue geometries
mesh # cells # dofs mortar # dofs !LM # dofs DG rmortar r!LM
@@
Figure 7: Comparison of system size for the mortar [18], multiplier-free [8] and the DG schemes in the case of interface dominated geometries. First order elements are used. Illustrations of interface dominated geometries with 5 (M1), 50 (M2), and 200 (M3) brain cells (ICS) surrounded by ECS (A). Comparison of the number of mesh cells (##\## cells), the number of degrees of freedom for the mortar scheme (##\## dofs mortar), the conforming multiplier-free scheme (##\## dofs !LM) and the DG scheme (##\## dofs DG) as well as the ratio between the system sizes of mortar and DG (rmortar) and of multiplier-free and DG (r!LM) for the three interface dominated meshes and the idealized 3D mesh from Model C (B).
@@ Acknowledgments
Ada J. Ellingsrud and Pietro Benedusi acknowledges support from the Research Council of Norway via FRIPRO grant #324239 (EMIx) and from the national infrastructure for computational science in Norway, Sigma2, via grant #NN8049K. Miroslav Kuchta acknowledges support from the Research Council of Norway via DataSim #303362. We thank Rami Masri for valuable discussions and input on discontinuous Galerkin methods and Jørgen Dokken for his contributions in setting up the associated software repository.